# EMOS: Embodiment-aware Heterogeneous Multi-robot Operating System with LLM Agents

### v1

@@ Abstract
Heterogeneous multi-robot systems (HMRS) have emerged as a powerful approach for tackling complex tasks that single robots cannot manage alone. Current large-language-model-based multi-agent systems (LLM-based MAS) have shown success in areas like software development and operating systems, but applying these systems to robot control presents unique challenges. In particular, the capabilities of each agent in a multi-robot system are inherently tied to the physical composition of the robots, rather than predefined roles. To address this issue, we introduce a novel multi-agent framework designed to enable effective collaboration among heterogeneous robots with varying embodiments and capabilities, along with a new benchmark named Habitat-MAS. One of our key designs is Robot Resume: Instead of adopting human-designed role play, we propose a self-prompted approach, where agents comprehend robot URDF files and call robot kinematics tools to generate descriptions of their physics capabilities to guide their behavior in task planning and action execution. The Habitat-MAS benchmark is designed to assess how a multi-agent framework handles tasks that require embodiment-aware reasoning, which includes 1) manipulation, 2) perception, 3) navigation, and 4) comprehensive multi-floor object rearrangement. The experimental results indicate that the robot’s resume and the hierarchical design of our multi-agent system are essential for the effective operation of the heterogeneous multi-robot system within this intricate problem context.
@@ 2 Related work
Multi-agent systems (MAS) have been a research focus for several decades. These systems consist of multiple interacting agents, which can be both cooperative and competitive, and are designed to solve problems that are difficult or impossible for a single agent to tackle alone Stone & Veloso (2000). The integration of LLMs into MAS is a relatively new yet rapidly growing area of research. This integration leverages the language understanding and generation capabilities of LLMs to enhance communication, coordination, and decision-making within MAS.  Wu et al. (2023a); Hong et al. (2023); Li et al. (2024) focus on the communication issues in LLM-based Multi-Agent Systems. Xu et al. (2024) proposes Crab, a cross-environment benchmark framework for evaluating Multimodal Language Models (MLMs) in different GUIs like mobile phones and desktop computers. For robotic intelligence,  Zhang et al. (2023) investigates how two agents can use communication to better collaborate and complete tasks in a multi-room scenario. Mandi et al. proposed RoCo, which is a multi-agent system for multi-arm collaboration. They try to leverage the 3D spatial reasoning capabilities to help multi-arm low-level trajectory planning. In comparison, we focus a more general mutli-agent scenario with drones, legged robots, wheeled robots with arms, with a multi-agent system required to understand general capabilities including navigation, manipulation and perception based on physics design.
MULTI-ROBOT SYSTEM
@@
Early works by Arai et al. (2002) and Ota (2006) laid the research foundation for multi-robot systems by providing a comprehensive overview of the progress and key challenges in MRS in around 2000, including MRS architectural design, distributed mapping, and navigation coordination, etc. Rizk et al. (2019) specifically reviewd the challenges in cooperative heterogeneous MRS, decomposing the MRS workflow to task decomposition, coalition formation, task allocation, perception and MRS planning and control. In this work, we also follow the established concept definitions and the principles of system design in this survey. Roldán et al. (2016) built a HMRS composed of aerial vehicles (drones) and ground vehicles to collaborate to monitor environmental variables of greenhouses. Kiener & Von Stryk (2010) designs a system composed of wheeled robot and humanoid robot to collaborate in a ”robot soccer” scenario. The authors carefully decompose the complex task into subtasks based on the robots’ capabilities, followed by human-crafted task allocation and planning algorithms. In this work, we focus on MRS for complex household environments with multiple floors, composed of drones, wheeled robots, and legged robots with robotic arms.
TASK PLANNING WITH LARGE LANGUAGE MODELS
@@
Large language models(LLMs) trained on massive corpora are generally considered to have acquired common sense knowledge for task planning (Vemprala et al., 2023; Yao et al., 2022; Zhao et al., 2023; Hu et al., 2023; Wu et al., 2024; Sha et al., 2023; Mu et al., 2024a; Gao et al., 2024). Thanks to recent advancements, directly generating plans with LLMs has become an active research area in recent years (Logeswaran et al., 2022; Wu et al., 2023b; Lin et al., 2023). When using LLMs for task planning, some approaches directly generate the entire plan in an open-loop manner, that is, without executing it in the environment (Huang et al., 2022a; Mu et al., 2023; Singh et al., 2022). An alternative line of research investigates closed-loop task planning, which offers greater flexibility for error correction, human interaction, and grounding the plan in the actual environmental state (Ahn et al., 2022; Guo et al., 2023; Huang et al., 2023; Hu et al., 2023; Huang et al., 2022b; Song et al., 2023; Hu et al., 2024; Mu et al., 2024b). This paper explores closed-loop task planning, where real-time environmental changes are integrated, and a central large language model processes these real-time changes and adapts plans accordingly.
@@ 3.4 Hierarchical Task Planning, Assignment and Action
To adapt the LLM-based MAS for real-time HMRS operation, the multi-agent action policy needs to be asynchronous, due to the potential asynchrony in multi-robot action execution. For this purpose, we design a hierarchical pipeline to perform task planning, assignment, and action execution for our LLM-based MAS. Specifically, there are two stages: 1) The first stage of Centralized Group Discussion runs in a synchronized fashion, in which all agents wait for messages from other agents, and the discussion history is seen by all agents. 2) While in the second stage of Decentralized Action Execution, each agent generates an action, waits for its execution it in the world, and generates a new action so on so forth. Each robot is associated with a robot-dedicated agent with full access to its robot resume to assist in decision-making and action execution. The pseudocode in Algorithm 1 provides a comprehensive overview of the hierarchical task planning, assignment, and then action execution within the EMOS framework.
In the first stage of centralized group discussion, there is a CentralPlanner that generates an initial plan for each robot, and each robot also has an LLM agent that checks its assigned subtask and provides feedback to the central planner in Reflection. With the robot resume composed of general description and numerical details, the robot-dedicated agent could reason its general availability for the assigned task and further check the geometric availability by using mathematical tools. Specifically, by accessing robot resumes in the last section, a robot-dedicated agent is encouraged to perform common-sense reasoning with textual summaries, and generate code with numerical data in both the scene description and the robot resume to do calculations for spatial-aware reasoning.
@@ 17            $history_{i}\leftarrow[history_{i},action_{i},response_{i}]$
In the second stage, with the result of the assignment of tasks, each robot-dedicated agent starts to execute its action in parallel. Given a subtask description and action execution history, a robot-dedicated agent controls the current agent by LLM FunctionCall with robot control libraries. These robot control libraries are implemented with ground truth world information, classic robot trajectory planners and inverse kinematics solvers. An agent automatically goes to wait state when it finishes all reasoned actions. When the agent fails to accomplish a subtask but ”believes” to have finished it, it could continue to execute the following planned actions. A side note is that, to study this circumstance of partial failure, we also evalute the task by sub-goals, which will be discussed in the experiments. The agent in the wait state will awake when it receives a new task assignment from a group discussion. For an episode, the task ends when all robots are in the wait state.
@@ 4.1 Benchmark Overview
The benchmark provides well-defined APIs for both task planning and robot control. For task planning and assignment, agents have access to tools for robot resume access construction and a Python code interpreter to execute the code during the first stage. Besides, for the robot control, robot agents have access to APIs that link to low-level robot skills, including fundamental actions such as navigate_to, move_arm_to, pick, and place, etc. Temporarily, these low-level robot skills are implemented with a classic robot trajectory planner and inverse kinematics solver. There are certain limitations in the skill libraries, such as the absence of explicit gripper control, which are addressed by contact-based grasping. As long as the robot gripper contacts an object, we snap the object to the gripper. One reason for this design is that, if we enable physics simulation for object-gripper contact simulation, it will introduce more irrelevant failures depending on the parameter tuning of physics simulation integration. Thus, we disable the integrated Pybullet (Coumans & Bai (2016–2021)) physics simulation in the benchmark. However, for tasks that require more sophisticated low-level robot control, the benchmark can be easily extended with re-enabling the physics simulation and robot learning-based policies.
@@ 4.2 Task Overview
There are four tasks carfully designed in the Habitat-MAS benchmark. Tasks 1, 2, and 3 aim to evaluate if agents are able to understand the three aforementioned robot capabilities respectively. Specifically, 1) Task 1 is deigned as a cross-floor object navigation task including two robots (wheeled and legged) navigating in a multi-floor scene, aiming to evaluate agent’s ability to understand robot’s mobility; 2) Task 2, named cooperative perception for manipulation, represents a common scenario in multi-robot collaboration, where robot perception assist manipulation. This task is set up to test the ability of MAS to reason about robot sensor type or view point in other word; 3) Task 3 is a classic household rearrangement task including two robots with different manipulation capabilities collaborate to manipulate object placed on specific receptacles, this task can cleverly test MAS’s ablility to understand robot arm’s workspace; 4) Task 4 is a multi-floor multi-agent and multi-object rearrangement task that requires the LLM-based multi-agent systems to comprehend all information and capabilities properly to collaborate. It is important to note that, during the creation of the benchmark dataset, we carefully filter the task episodes so that each robot in the scene can only complete a subset of the subgoals. In other words, the multi-agent system must comprehend robots’ physical capabilities to forge a feasible plan. For detailed task description, refer to Appendix A.2.
@@ 4.3 Evaluation Criteria
The performance of multi-agent collaboration in Habitat-MAS is evaluated using several key metrics: 1) Success Rate. Based on the task design we introduced in the last section, we define a series of intermediate subgoals in PDDL language for each task to evaluate the task result. This metric evaluates the proportion of episodes in which an MRS successfully completes all sub-goals, which directly reflects the overall planning and coordination capabilities of MAS. 2) Sub-goal Success Rate. This metric calculates the percentage of sub-goals achieved by the MRS. Due to the limit on pages, please refer to Appendix A.3.1 for more details about the sub-goal definition and implementation. 3) Token Usage. The used tokens are a key metric to evaluate the efficiency of LLM-based MAS. The effectiveness of agent communication and action planning is measured by the number of tokens used during discussions. This reflects how efficiently agents coordinate and strategize to complete tasks. 4) Simulation Step. We also evaluate the number of simulation steps consumed by the MAS to complete each task. Drones typically move the fastest, followed by wheeled robots, with legged robots being the slowest. This metric evaluates the LLM-based MAS’ ability to assign tasks for high MRS efficiency. For instance, in an extreme scenario, one robot handles all the subtasks, leaving the rest of the robots without any assignments. This situation results in low efficiency for the HMRS and causes an abnormally large number of simulation steps.
@@ 4.4 Experiments with EMOS
In this section, we present the experimental result of our EMOS system on our Habitat-MAS benchmark, along with ablation studies to explain the impact of different building blocks. Our benchmark offers a large-scale dataset with episodes in more than 70 distinct scenes. However, due to budget constraints, all ablation studies were conducted on a subset of 519 episodes. We use the GPT-4o (OpenAI, 2024) API of the May 2024 version in this experiment. For more details on how episodes are generated and the full set of those episodes, please refer to Appendix A.3.2.
@@
We present the numerical results of the ablation study in Table 1. Firstly, by comparing the EMOS with EMOS w/o. numerical description, we observe that LLM agents can still perform relatively well in simple tasks like navigation and perception, as in Figure 5. We infer this is because LLM agents can still understand robots’ mobility from the tree structure of URDF, recognizing the node names like wheel, leg, etc. Additionally, LLM agents can infer from their common sense that Drone is an aerial robot with a relatively broad view without any explicit information about Drone’s camera height in the input URDF. However, the success rate drops significantly in more complex tasks like single-floor rearrangement (28.35% $\rightarrow$ 9.20%) and multi-floor rearrangement (13.46% $\rightarrow$ 3.85%). This emphasizes that mere textual descriptions are inadequate for robotic tasks requiring precise manipulation. In this case, invoking mathematical functions to process and reflect numerically helps. Secondly, in the setting w/o. robot resume, by further removing the textual summary extracted from URDF, the success rate in navigation (37.37% $\rightarrow$ 14.14%) and perception (52.94% $\rightarrow$ 28.32%) tasks both decrease dramatically in the experiments on MAS without robot resume. This result proves that LLM agents can indeed be aware of robot embodiment capabilities through commonsense reasoning, rather than human-assigned role-playing. By comparing the average task success rate of this ablated setting with our intact EMOS system (37.82% $\rightarrow$ 15.63%), as in Table 1, it confirms the superiority of our methods utilizing both numerical reasoning with tools and textual reasoning with common sense. Thirdly, compared to EMOS, the setting w/o. group discussion performs the worst on success rate (37.82% $\rightarrow$ 15.23%) as demonstrated in Table 1. In this setting, each robot agent executes tasks directly according to task description, without leader assignments or self-reflection on their embodiedment limitation using mathematical tools. Although this setting significantly reduces token usage by a large margin compared to the other three methods, it also dramatically reduces the success rate. In addition, when inspecting the column of simulation steps in Table 1, EMOS can complete the tasks in the second least steps on average, while EMOS w/o. robot resume struggles in planning, consuming the most steps, and EMOS w/o. group discussion ends the episodes with the least steps due to failure. For special case study, please refer to Appendix A.4. For more detailed discussions on experiments, see Appendix A.5.

### v2

@@ Abstract
    Heterogeneous multi-robot systems (HMRS) have emerged as a powerful approach for tackling complex tasks that single robots cannot manage alone. Current large-language-model-based multi-agent systems (LLM-based MAS) have shown success in areas like software development and operating systems, but applying these systems to robot control presents unique challenges. In particular, the capabilities of each agent in a multi-robot system are inherently tied to the physical composition of the robots, rather than predefined roles. To address this issue, we introduce a novel multi-agent framework designed to enable effective collaboration among heterogeneous robots with varying embodiments and capabilities, along with a new benchmark named Habitat-MAS. One of our key designs is Robot Resume: Instead of adopting human-designed role play, we propose a self-prompted approach, where agents comprehend robot URDF files and call robot kinematics tools to generate descriptions of their physics capabilities to guide their behavior in task planning and action execution. The Habitat-MAS benchmark is designed to assess how a multi-agent framework handles tasks that require embodiment-aware reasoning, which includes 1) manipulation, 2) perception, 3) navigation, and 4) comprehensive multi-floor object rearrangement. The experimental results indicate that the robot’s resume and the hierarchical design of our multi-agent system are essential for the effective operation of the heterogeneous multi-robot system within this problem context. The project website is: https://emos-project.github.io/
@@ 2 Related work
    LLM-Based Multi-Agent System. The integration of LLMs into MAS is a relatively new yet rapidly growing area of research. This integration leverages the language understanding and generation capabilities of LLMs to enhance communication, coordination, and decision-making within MAS.  Wu et al. (2023a); Hong et al. (2023); Li et al. (2024) focus on the communication issues in LLM-based Multi-Agent Systems. Xu et al. (2024) proposes Crab, a cross-environment benchmark framework for evaluating Multimodal Language Models (MLMs) in different GUIs like mobile phones and desktop computers. For robotic intelligence,  Zhang et al. (2023) investigates how two agents can use communication to better collaborate and complete tasks in a multi-room scenario. Mandi et al. proposed RoCo, which is a multi-agent system for multi-arm collaboration. They try to leverage the 3D spatial reasoning capabilities to help multi-arm low-level trajectory planning. In comparison, we focus a more general mutli-agent scenario with drones, legged robots, wheeled robots with arms, with a multi-agent system required to understand general capabilities including navigation, manipulation and perception based on physics design.
    Heterogeneous Multi-Agent Learning. Heterogeneous multi-agent systems involve agents with varying capabilities or functional roles working collaboratively toward shared goals. This field has gained significant attention due to its practical relevance in real-world applications requiring diverse agent teams. Recent works have introduced innovative frameworks for learning and coordination in heterogeneous teams. For example, Seraj et al. (2022) proposes a method for learning communication protocols tailored to each robot’s role and capabilities, optimizing team performance in dynamic environments. Similarly, Bettini et al. (2023) develops reinforcement learning algorithms specifically designed for heterogeneous teams, enabling effective inter-agent coordination despite differences in robot traits. Task allocation is another critical aspect. Ravichandar et al. (2020) provides a scalable optimization framework for balancing workload and resource utilization across large heterogeneous teams. In addition, Seraj et al. (2024b) combines human demonstrations with machine learning to train diverse robot teams efficiently. Seraj et al. (2024a) introduces a novel policy network architecture that integrates individual robot policies into a composite framework for effective decision-making. Our work distinguishes these works in the way that we handle the behavior of these heterogeneous agents by transfering the prior knowledge in the pre-trained large-language models without extra training.
@@
    Multi-Robot System. Early works by Arai et al. (2002) and Ota (2006) laid the research foundation for multi-robot systems by providing a comprehensive overview of the progress and key challenges in MRS in around 2000, including MRS architectural design, distributed mapping, and navigation coordination, etc. Hamann & Wörn (2008) proposed a model framework with an explicit space representation for swarm robotic algorithm design, deriving an abstract swarm motion model from a single robot description and validating it against simulation results, while also discussing the challenges and related work in this area. Rizk et al. (2019) specifically reviewd the challenges in cooperative heterogeneous MRS, decomposing the MRS workflow to task decomposition, coalition formation, task allocation, perception and MRS planning and control. In this work, we also follow the established concept definitions and the principles of system design in this survey. Roldán et al. (2016) built a HMRS composed of aerial vehicles (drones) and ground vehicles to collaborate to monitor environmental variables of greenhouses. Kiener & Von Stryk (2010) designs a system composed of wheeled robot and humanoid robot to collaborate in a ”robot soccer” scenario. The authors carefully decompose the complex task into subtasks based on the robots’ capabilities, followed by human-crafted task allocation and planning algorithms. Yang & Parasuraman (2020) proposed Self-Adaptive Swarm System (SASS), a hierarchical needs-based framework for cooperative multi-robot systems, inspired by Maslow’s hierarchy of human needs, combining multi-robot capabilities with a distributed negotiation-agreement mechanism that prioritizes robots’ needs according to Maslow’s human needs principle.
@@
    Task Planning With Large Language Models. Large language models(LLMs) trained on massive corpora are generally considered to have acquired common sense knowledge for task planning (Vemprala et al., 2023; Yao et al., 2022; Zhao et al., 2023). Thanks to recent advancements, directly generating plans with LLMs has become an active research area in recent years (Logeswaran et al., 2022; Wu et al., 2023b; Lin et al., 2023). When using LLMs for task planning, some approaches directly generate the entire plan in an open-loop manner, that is, without executing it in the environment (Huang et al., 2022a; Mu et al., 2023; Singh et al., 2022). An alternative line of research investigates closed-loop task planning, which offers greater flexibility for error correction, human interaction, and grounding the plan in the actual environmental state (Ahn et al., 2022; Guo et al., 2023; Huang et al., 2023; Hu et al., 2023; Huang et al., 2022b; Song et al., 2023; Hu et al., 2024). This paper explores closed-loop task planning, where real-time environmental changes are integrated, and a central large language model processes these real-time changes and adapts plans accordingly.
@@ 3.4 Hierarchical Task Planning, Assignment and Action
    To adapt the LLM-based MAS for real-time HMRS operation, the multi-agent action policy needs to be asynchronous, due to the potential asynchrony in multi-robot action execution. For this purpose, we design a hierarchical pipeline to perform task planning, assignment, and action execution for our LLM-based MAS. Specifically, there are two stages: 1) The first stage of Centralized Group Discussion runs in a synchronized fashion, in which all agents wait for messages from other agents, and the discussion history is seen by all agents. 2) While in the second stage of Decentralized Action Execution, each agent generates an action, waits for its execution it in the world, and generates a new action so on so forth. Each robot is associated with a robot-dedicated agent with full access to its robot resume to assist in decision-making and action execution. The pseudocode in Algorithm 1 provides a comprehensive overview of the hierarchical task planning, assignment, and then action execution within the EMOS framework. For detailed explanation for Algorithm 1 and MAS design, please refer to Appendix A.2
@@ 4.1 Benchmark Overview
    The benchmark provides well-defined APIs for both task planning and robot control. For task planning and assignment, agents have access to tools for robot resume access construction and a Python code interpreter to execute the code during the first stage. Besides, for the robot control, robot agents have access to APIs that link to low-level robot skills, including fundamental actions such as navigate_to, move_arm_to, pick, and place, etc. Temporarily, these low-level robot skills are implemented with a classic robot trajectory planner and inverse kinematics solver. There are certain limitations in the skill libraries, such as the absence of explicit gripper control, which are addressed by contact-based grasping. As long as the robot gripper contacts an object, we snap the object to the gripper. One reason for this design is that, if we enable physics simulation for object-gripper contact simulation, it will introduce more irrelevant failures depending on the parameter tuning of physics simulation integration. Thus, we disable the integrated Pybullet (Coumans & Bai (2016–2021)) physics simulation in the benchmark. However, for tasks that require more sophisticated low-level robot control, the benchmark can be easily extended with re-enabling the physics simulation and robot learning-based policies. For more detailed explaination about simulation and robot low-level control, please refer to Appendix B.4.
@@ 4.2 Task Overview
    There are four tasks carfully designed in the Habitat-MAS benchmark. Tasks 1, 2, and 3 aim to evaluate if agents are able to understand the three aforementioned robot capabilities respectively. Specifically, 1) Task 1 is deigned as a cross-floor object navigation task including two robots (wheeled and legged) navigating in a multi-floor scene, aiming to evaluate agent’s ability to understand robot’s mobility; 2) Task 2, named cooperative perception for manipulation, represents a common scenario in multi-robot collaboration, where robot perception assist manipulation. This task is set up to test the ability of MAS to reason about robot sensor type or view point in other word; 3) Task 3 is a classic household rearrangement task including two robots with different manipulation capabilities collaborate to manipulate object placed on specific receptacles, this task can cleverly test MAS’s ablility to understand robot arm’s workspace; 4) Task 4 is a multi-floor multi-agent and multi-object rearrangement task that requires the LLM-based multi-agent systems to comprehend all information and capabilities properly to collaborate. It is important to note that, during the creation of the benchmark dataset, we carefully filter the task episodes so that each robot in the scene can only complete a subset of the subgoals. In other words, the multi-agent system must comprehend robots’ physical capabilities to forge a feasible plan. For detailed task description, refer to Appendix B.1.
@@ 4.3 Evaluation Criteria
    The performance of multi-agent collaboration in Habitat-MAS is evaluated using several key metrics: 1) Success Rate. Based on the task design we introduced in the last section, we define a series of intermediate subgoals in PDDL language for each task to evaluate the task result. This metric evaluates the proportion of episodes in which an MRS successfully completes all sub-goals, which directly reflects the overall planning and coordination capabilities of MAS. 2) Sub-goal Success Rate. This metric calculates the percentage of sub-goals achieved by the MRS. Due to the limit on pages, please refer to Appendix B.2 for more details about the sub-goal definition and implementation. 3) Token Usage. The used tokens are a key metric to evaluate the efficiency of LLM-based MAS. The effectiveness of agent communication and action planning is measured by the number of tokens used during discussions. This reflects how efficiently agents coordinate and strategize to complete tasks. 4) Simulation Step. We also evaluate the number of simulation steps consumed by the MAS to complete each task. Drones typically move the fastest, followed by wheeled robots, with legged robots being the slowest. This metric evaluates the LLM-based MAS’ ability to assign tasks for high MRS efficiency. For instance, in an extreme scenario, one robot handles all the subtasks, leaving the rest of the robots without any assignments. This situation results in low efficiency for the HMRS and causes an abnormally large number of simulation steps.
@@ 4.4 Experiments with EMOS
    In this section, we present the experimental result of our EMOS system on our Habitat-MAS benchmark, along with ablation studies to explain the impact of different building blocks. Our benchmark offers a large-scale dataset with episodes in more than 70 distinct scenes. However, due to budget constraints, all ablation studies were conducted on a subset of 519 episodes. We use the GPT-4o (OpenAI, 2024) API of the May 2024 version in this experiment. For more details on how episodes are generated and the full set of those episodes, please refer to Appendix B.3.
@@
    We present the numerical results of the ablation study in Table 1. Firstly, by comparing the EMOS with EMOS w/o. numerical description, we observe that LLM agents can still perform relatively well in simple tasks like navigation and perception, as in Figure 5. We infer this is because LLM agents can still understand robots’ mobility from the tree structure of URDF, recognizing the node names like wheel, leg, etc. Additionally, LLM agents can infer from their common sense that Drone is an aerial robot with a relatively broad view without any explicit information about Drone’s camera height in the input URDF. However, the success rate drops significantly in more complex tasks like single-floor rearrangement (28.35% $\rightarrow$ 9.20%) and multi-floor rearrangement (13.46% $\rightarrow$ 3.85%). This emphasizes that mere textual descriptions are inadequate for robotic tasks requiring precise manipulation. In this case, invoking mathematical functions to process and reflect numerically helps. Secondly, in the setting w/o. robot resume, by further removing the textual summary extracted from URDF, the success rate in navigation (37.37% $\rightarrow$ 14.14%) and perception (52.94% $\rightarrow$ 28.32%) tasks both decrease dramatically in the experiments on MAS without robot resume. This result proves that LLM agents can indeed be aware of robot embodiment capabilities through commonsense reasoning, rather than human-assigned role-playing. By comparing the average task success rate of this ablated setting with our intact EMOS system (37.82% $\rightarrow$ 15.63%), as in Table 1, it confirms the superiority of our methods utilizing both numerical reasoning with tools and textual reasoning with common sense. Thirdly, compared to EMOS, the setting w/o. group discussion performs the worst on success rate (37.82% $\rightarrow$ 15.23%) as demonstrated in Table 1. In this setting, each robot agent executes tasks directly according to task description, without leader assignments or self-reflection on their embodiedment limitation using mathematical tools. Although this setting significantly reduces token usage by a large margin compared to the other three methods, it also dramatically reduces the success rate. In addition, when inspecting the column of simulation steps in Table 1, EMOS can complete the tasks in the second least steps on average, while EMOS w/o. robot resume struggles in planning, consuming the most steps, and EMOS w/o. group discussion ends the episodes with the least steps due to failure. For special case study, please refer to Appendix C.1. For more detailed discussions on experiments, see Appendix C.